# Galley: Modern Query Optimization for Sparse Tensor Programs

### v4

@@ 5.3. Variable Elimination
The core of our logical optimizer is the variable elimination (VE) framework. With this view, the logical plan for an input query is defined by an order on the indices being aggregated over, i.e., an elimination order. Given this order, we can construct a valid logical plan by iterating through the elimination order one index at a time in order to (1) identify the minimal sub-expression needed to aggregate over it, (2) create a new logical query representing the result of that sub-expression, and (3) replace it in the original query with an alias to the result. At the end of this process, the remaining query no longer requires any aggregation and therefore is itself a logical query.
@@
The elimination order $jkl$ corresponds to a left-to-right multiplication strategy because eliminating $j$ from the expression first requires performing the matrix multiplication between $A$ and $B$. Eliminating $k$ then requires multiplying that intermediate result with $C$, and so on. Concretely, this produces the following sequence of logical queries:
@@
Similarly, the elimination order $lkj$ corresponds to a right-to-left strategy, and the order $klj$ to a middle-first strategy.
@@ 5.4. Identifying Minimal Sub-Expressions
We now explain how to identify the minimal sub-expressions (MSEs) needed to eliminate an index. In sum-product expressions, the MSE is simply a summation over that index wrapping the product of the tensors that are indexed by it. However, it is not immediately clear how to do this for more complex input programs. Fortunately, we show that identifying MSEs corresponds to a careful traversal down the annotated expression tree, examining the algebraic properties of the operation at each node to determine how to proceed.
Figure 5. Annotated expression tree for logistic regression
$\sigma(\sum_{jpc}S_{ipc}(P_{pj}\theta_{j}+C_{cj}\theta_{j}))$
@@
Fig. 5 shows the annotated expression tree for our logistic regression example after distributing the multiplication as in Eq. (2). The sigmoid function is the outermost layer of the expression, so it appears at the top of the tree. The summations all occur just inside the sigmoid function, so they annotate the top multiplication operator, as denoted by the red arrow.
@@
Distributive Functions. When we reach a function that distributes over the aggregate (e.g., * and $\sum$), we examine how many of the children, subtrees of the AET, contain the current index. If one child contains the index, we traverse down that child’s branch, i.e., we factor the other children out of the aggregate. If multiple children contain the index, we cannot go any further down the tree, so we wrap the sub-tree rooted at that node in the aggregate and return it as our MSE. If the function is commutative and associative, we are slightly more precise and include only the children that contain the index.
Commutative, Identical Functions. When the node’s function is the same as the aggregate function and is commutative, we can push the aggregate down to each child independently. For example, we can transform the expression $\sum_{i}A_{i}+B_{i}$ into $\sum_{i}A_{i}+\sum_{i}B_{i}$. For all children that contain the index, we add the result of traversing down its branch to the list of MSEs and replace it with an alias to the result. If a child does not contain the index, then we need to account for the repeated application of the aggregate function. To do this, we require a function $g(x,n)=f(x,...,x)$ that represents the repeated application of our aggregate function $f$. When $f$ is addition, $g$ is multiplication. When $f$ is idempotent, $g(x,n)=x$. We then wrap each non-index child in a Map with function $g$ and the size of the index’s dimension as the second child. Further, we can choose to push down along some of these dimensions and not others, e.g. $\sum_{i}(A_{i}+B_{i}+C_{i})=\sum_{i}(A_{i}+B_{i})+\sum_{i}(C_{i})$. To allow this, we represent each path that a variable can be pushed down with a virtual index, and we include these in our elimination ordering. This increases our optimization space but allows us to capture more aggregation strategies.
@@
Discussion. Galley builds upon and extends the FAQ framework for optimizing conjunctive queries with aggregation(Khamis et al., 2016). This framework explored the optimization of queries with the following form, where each $\bigoplus^{(i)}$ is either equal to or forms a semi-ring with $\otimes$:
@@
Though this framework captures many important problems, it lacks the flexibility needed to support a general tensor processing system. Consider a slightly modified version of the SDDMM kernel:
@@
This expression is not an FAQ query because it mixes addition and multiplication in the pointwise expression. Similarly, our logistic regression example, $\sigma(\sum_{j}X_{ij}\theta_{j})$, cannot be expressed as an FAQ because the aggregate occurs within $\sigma$. Galley extends this framework by accommodating arbitrary pointwise function composition and arbitrary placement of aggregates within expressions.
@@ 5.5. Restricted Elimination Orders
Before we can search for the optimal elimination order, we must define the space of valid elimination orders. Depending on the input’s structure, the order in which indices can be eliminated might be restricted. This could be due to non-commutative aggregates or aggregate placement. The former occurs when an aggregate wraps another aggregate it does not commute with. For example, if the expression is $\max_{i}\sum_{j}A_{ij}$, we must perform the summation before handling the maximum because max and $\sum$ do not commute. The placement issue arises when an aggregate wraps another aggregate but cannot reach it via the traversal described above, e.g., $\sum_{i}\sqrt{\sum_{j}A_{ij}}$; in this case, the inner aggregate must be performed first. Collectively, these restrictions form a partial ordering on the index variables that must be respected when we enumerate elimination orders.
@@ 5.6. Search Algorithms
The elimination order for this expression is an ordering of the indices $\{p,c,j^{1},j^{2}\}$. Note that $j^{1}$ and $j^{2}$ are virtual indices, introduced to represent the two paths that $j$ can be pushed down in the AET due to the addition operator. Galley’s logical optimizer searches through these possible orders to find the most efficient one. In this case, it would choose $[j^{1},j^{2},p,c]$, resulting in the following logical plan,
@@
Greedy. The greedy approach simply chooses the cheapest index to aggregate at each point by finding the minimal sub-query for each index and computing its cost. The cheapest index’s minimal sub-query is removed from the expression and appended to the logical plan; it is then replaced in the remaining query with an alias to the result. This continues until no aggregates remain in the expression.
Branch-and-Bound. The branch-and-bound approach computes the optimal variable order and occurs in two steps. The first step uses the greedy algorithm to produce an upper bound on the cost of the overall plan; the second performs a dynamic programming algorithm. In the dynamic programming step, the keys of the memo table are unordered sets of indices, and the values are a tuple of partial elimination orders, residual queries, and costs. The algorithm initializes the table with the empty set and a cost of zero. At each step, it iterates through table entries and attempts to aggregate out another index. It then uses the cost bound from the first step to prune entries from the memo table whose cost exceeds the bound; doing so is valid because costs monotonically increase as more indices are added to the set. At the end of this step, the algorithm returns the index order associated with the full set of indices.
@@ 6. Physical Optimizer
Each query in the logical dialect roughly corresponds to a single loop nest and materialized intermediate. However, several decisions remain about how the kernel is computed, including: (1) the loop order over the indices, (2) the format (i.e., layout) of the result, and (3) the protocol (i.e., search algorithm) for accessing each index of each input. The physical optimizer makes these decisions.
@@ 6.1. Loop Order
The loop order of a tensor kernel determines the order of access to the inputs. An optimal order results in early pruning of iterations due to early intersection of sparse inputs. Intuitively, this is similar to selecting a variable order for a worst-case optimal join algorithm. To find the optimal order, Galley has a cost model for loop orders based on sparsity estimates and an optimization algorithm that uses this cost model to guide its search.
@@
The cost of a loop order equals the number of iterations each level of the loop nest incurs.
@@
Suppose that $A$ has only a single non-zero entry and that $B$ and $C$ have 5 non-zero entries per column and per row. In this case, the loop order $ijkl$ is significantly more efficient than $lkji$. In the former, the first two loops, over $i$ and $j$, incur only a single iteration because they are bounded by the size of $A$. The third and fourth incur $5$ and $5^{2}$ iterations, respectively, because there are only $5$ non-zero $k$’s per $j$ in $B$ and $5$ non-zero $l$’s per $k$ in $C$. In the latter, the first two loops iterate over the full matrix $C$ despite most of those iterations not leading to useful computation.
The last piece of our cost model is the cost of transposition. If an input’s index order is not concordant with the loop order, it must be transposed before the query can be executed. This imposes a cost that is linear in that input’s size. We ignore the cost of transposing intermediates because we they are always materialized in a concordant order.
Formally, let $Q$ be the pointwise expression in our kernel, and let $Q_{(i_{1},\ldots,i_{k})}$ be the restriction of that expression to just the sequence of index variables $i_{1},\ldots,i_{k}$. Further, let $\mathbf{A}^{(i_{1},\ldots,i_{k})}$ be the set of input tensors which are not concordant with $i_{1},\ldots,i_{k}$. Then, we can define the cost of a loop order as follows,
@@
In practice, we further refine this model to take into account the number and kind of tensor accesses at each level of the loop nest.
@@
To optimize the loop order, we combine this cost model with a branch-and-bound, dynamic programming algorithm. In the first pass, the optimization algorithm selects the cheapest loop index at each step until reaching a full loop order. This produces an upper bound on the optimal execution cost, which the algorithm uses to prune loop orders in the second step. This step applies a dynamic programming algorithm. Taking inspiration from Selinger’s algorithm for join ordering, each key in the DP table is a set of index variables and a set of inputs. The former represent the loops that have been iterated so far, and the latter represents a set of inputs that must be transposed.
@@ 6.2. Intermediate Formats
Once the loop order has been determined, the physical optimizer selects the optimal format for each query’s output. First, Galley selects the order of the indices in the output to be concordant with either the loop order of the kernel where it will be consumed or the order requested by the user. It then selects a format for each index of the output (e.g., dense vector, hash table, etc.). Two factors affect this decision: (1) the kind of writes being performed (sequential vs random) and (2) the sparsity of the tensor at this index. The former is important because many formats (e.g., sorted list formats) permit only sequential construction. These formats can only be used if the output indices up to that point form a prefix of the loop order.
When considering sparsity, Galley balances the fact that denser formats tend to be more efficient due to their memory locality and simplicity, while sparser formats are asymptotically smaller for sparse outputs. To describe this trade-off, we hand selected sparsity cutoffs between fully sparse, bytemap, and fully dense formats. To determine a particular output index’s format, the physical optimizer first determines the sparsity at this index level and uses our cutoffs to determine which category of formats to consider. Then, it checks whether sequential or random writes are being performed and selects the most efficient format that supports the write pattern.
@@ 7.1. Sparsity and Cardinality Estimation
Sparsity estimation is highly related to cardinality estimation in databases. However, translating methods for the latter to the former requires analyzing the algebraic properties of our tensor programs. For example, let $A_{ij}$ and $B_{jk}$ be sparse matrices with a fill value of 00, and let $R_{A}(I,J)$ and $R_{B}(J,K)$ be relations that store the indices of their non-zero entries. Assume we are performing the following,
@@
The correspondence results from the fact that 00 is the annihilator of multiplication (i.e., $x*0=0\forall x$), so any non-zero entry $ijk$ in the output must correspond to a non-zero $ij$ in $A$ and a non-zero $jk$ in $B$. Consider the following instead:
@@
In this case, a nonzero $ijk$ in the output can result from a non-zero $ij$ in $A$ or a non-zero $jk$ in $B$. In traditional relational algebra, where relations are over infinite domains, this kind of disjunction would result in an infinite relation. However, tensors have finite dimensions, so we can introduce relations that represent the finite domains of each index, e.g., $D_{i}=\{1,...,n_{i}\}$. This lets us represent the index relation of the output as
@@ 7.2. The Sparsity Statistics Interface
We use our statistics interface to annotate an expression with stats objects at every node of the AST in a bottom-up fashion. Each stats object then represents the sparsity patterns of the intermediate tensor output from that node. Surprisingly, to support sparsity estimation over the varied workloads and arbitrary operators of tensor algebra, we need to implement only a few core functions: (1) a constructor, which produces statistics from a materialized tensor for Input and Alias nodes, (2) a function for annihilating Map nodes (i.e., those whose children’s fill values are the annihilator of its pointwise function), which merges the children’s statistics, (3) a function for non-annihilating Map nodes, which merges the children’s statistics, (4) a function for Agg, which adjusts the input’s statistics to reflect an aggregation over some set of indices, and (5) an estimation procedure, which estimates the number of non-fill entries based on statistics about a tensor.
@@ 7.3.1. Uniform Estimator
The simplest statistic that can be kept about a tensor is the number of non-fill (e.g., non-zero) entries. The uniform estimator uses only this statistic and relies on the assumption that these entries are uniformly distributed across the dimension space. This corresponds to System-R’s cardinality estimator with the added assumption that the size of the index attribute’s active domain equals the size of the dimension (Selinger et al., 1979).
Constructor
Given a tensor $A_{i_{1},\ldots,i_{k}}\in\mathbf{R}^{n_{i_{1}}\times\ldots\times n_{i_{k}}}$, this function simply counts the non-fill values in the tensor, $nnz(A)$, and notes the dimension sizes $n_{i_{1}},\ldots,n_{i_{k}}$.
Map (Annihilating)
@@
Map (Non-Annihilating)
To handle an non-annihilating pointwise operation, this function calculates the probability that an entry in the output was fill in all inputs. Then, it takes the compliment to get the probability that it was non-fill in all inputs and multiplies this by the output dimension space. Using the preceding notation:
@@
Aggregate
Given an input tensor $A_{I}$ to aggregate over the indices $I^{\prime}$, this function computes the probability that an output entry is non-fill by calculating the probability that at least one entry in the subspace of the input tensor was not fill:
@@
Estimate
The estimation function simply returns the current tensor’s stored cardinality statistic.

### v5

@@ 4.3. Variable Elimination
The core of our logical optimizer is an extension of the variable elimination (VE) (eq. FAQ) framework (Dechter, 1999; Khamis et al., 2016). In its original context, this algorithm described a means of marginalization for probabilistic models by removing one variable at a time. When applied to our setting, it allows us to define the logical plan for an input query via an order on the indices being aggregated over, i.e., an elimination order. If we are given this order, we can construct a valid logical plan by iterating through the elimination order one index at a time in order to (1) identify the minimal sub-expression needed to aggregate over it, (2) create a new logical query representing the result of that sub-expression, and (3) replace it in the original query with an alias to the result. At the end of this process, the remaining query no longer requires any aggregation and therefore is itself a logical query.
@@
The elimination order jkljkl corresponds to a left-to-right multiplication strategy because eliminating $j$ from the expression first requires performing the matrix multiplication between $A$ and $B$. Eliminating $k$ then requires multiplying that intermediate result with $C$, and so on. Concretely, this produces the following sequence of logical queries:
@@
Similarly, the elimination order lkjlkj corresponds to a right-to-left strategy, and the order kljklj to a middle-first strategy.
@@ 4.4. Identifying Minimal Sub-Expressions
We now explain how to identify the minimal sub-expressions (MSEs) needed to eliminate an index. In sum-product expressions, the MSE is simply the product of the tensors that are indexed by it. For more complex input programs, we show that identifying MSEs corresponds to a careful traversal down the annotated expression tree, examining the algebraic properties of each operation to determine how to proceed.
Figure 5. Annotated expression tree for logistic regression over joins $\sigma(\sum_{jpc}S_{ipc}(P_{pj}\theta_{j}+C_{cj}\theta_{j}))$
@@
Fig. 5 shows the annotated expression tree for logistic regression where the input matrix is defined by a join-like expression $X_{ij}=S_{ipc}(P_{pj}+C_{cj})$. Further, Galley has pushed down $\theta_{j}$ into this expression. The sigmoid function is the outermost layer of the expression, so it appears at the top of the tree. The summations all occur just inside the sigmoid function, so they annotate the top multiplication operator.
@@
Distributive Functions. When we reach a function that distributes over the aggregate (e.g., * and $\sum$), we examine how many of the children, subtrees of the AET, contain the current index. If one child contains the index, we traverse down that child’s branch, i.e., we factor the other children out of the aggregate. If multiple children contain the index, we wrap the sub-tree rooted at that node in the aggregate and return it as our MSE. If the function is commutative and associative, we only include the children that contain the index.
Commutative, Identical Functions. When the node’s function is the same as the aggregate function and is commutative, we can push the aggregate down to each child independently. For example, we can transform the expression $\sum_{i}A_{i}+B_{i}$ into $\sum_{i}A_{i}+\sum_{i}B_{i}$. For all children that contain the index, we add the result of traversing down its branch to the list of MSEs and replace it with an alias to the result. If a child does not contain the index, then we need to account for the repeated application of the aggregate function. For example, $\sum_{i}B=N_{i}*B$ where $N_{i}$ is the size of the $i$ dimension.
@@
Discussion. Galley builds upon and extends the theoretical FAQ framework for optimizing conjunctive queries with aggregation(Khamis et al., 2016). This framework explored the optimization of queries with the following form, where each $\bigoplus^{(i)}$ is either equal to or forms a semi-ring with $\otimes$:
@@
Similarly to Galley, the FAQ paper described the optimization problem as selecting an optimal elimination order over the aggregated variables. Though this framework captures many important problems, it lacks the flexibility needed to support a general tensor processing system. Consider a slightly modified version of the SDDMM kernel:
@@
This expression is not an FAQ query because it mixes addition and multiplication in the pointwise expression. Galley extends this framework to accommodate arbitrary pointwise expressions and placement of aggregates within expressions.
@@ 4.5. Restricted Elimination Orders
Depending on the program structure, the order in which indices can be eliminated might be restricted. This could be due to non-commutative aggregates or aggregate placement. The former is when an aggregate wraps another aggregate that it does not commute with. For example, given $\max_{i}\sum_{j}A_{ij}$, we must perform the summation before handling the maximum because $\max$ and $\sum$ do not commute. The latter is when an aggregate wraps another aggregate but cannot reach it via the traversal described above, e.g., $\sum_{i}\sqrt{\sum_{j}A_{ij}}$; in this case, the inner aggregate must be performed first. Collectively, these restrictions form a partial ordering on the index variables that must be respected when we enumerate elimination orders.
@@ 4.6. Search Algorithms
The elimination order for this expression is an ordering of the indices $\{p,c,j\}$. Galley’s logical optimizer searches through these possible orders to find the most efficient one. In this case, it would choose $[j,p,c]$, resulting in the following logical plan,
@@
Greedy. The greedy approach chooses the cheapest index to aggregate at each step by finding the minimal sub-query for each index and computing its cost. The cheapest index’s minimal sub-query is removed from the expression, appended to the logical plan, and replaced with an alias to the result. This continues until no aggregates remain in the expression.
Branch-and-Bound. The branch-and-bound approach computes the optimal variable order and occurs in two steps. The first step uses the greedy algorithm to produce an upper bound on the cost of the overall plan; the second performs a dynamic programming algorithm. In the dynamic programming step, the keys of the memo table are unordered sets of indices, and the values are tuples containing a partial elimination order, residual query, and cost. The algorithm initializes the table with the empty set and a cost of zero. At each step, it iterates through table entries and attempts to aggregate out another index. It then uses the cost bound from the first step to prune entries from the memo table whose cost exceeds the bound; doing so is valid because costs monotonically increase as more indices are added to the set. At the end of this step, the algorithm returns the index order associated with the full set of indices.
@@ 5. Physical Optimizer
Each query in the logical dialect roughly corresponds to a single loop nest and materialized intermediate. However, several decisions remain about how the kernel is computed, including: (1) the loop order over the indices, (2) the format of the result, and (3) the merge algorithm for each index. The physical optimizer makes these decisions.
@@ 5.1. Loop Order
The loop order determines that inputs are accessed. An good loop order prunes the iteration space due to early intersection of sparse inputs. Intuitively, this is similar to selecting a variable order for a worst-case optimal join algorithm. Galley’s physical optimizer searches the space of loop orders to find one with the minimum cost, defined below.
@@
The cost of a loop order is composed of each loop’s number of iterations and the cost of transposing inputs to make them concordant with the loop order.
@@
Suppose that $A$ has only a single non-zero entry and that $B$ and $C$ have 5 non-zero entries per column and per row. In this case, the loop order ijklijkl is significantly more efficient than lkjilkji. In the former, the first two loops, over $i$ and $j$, incur only a single iteration because they are bounded by the size of $A$. The third and fourth incur $5$ and $5^{2}$ iterations, respectively, because there are only $5$ non-zero $k$’s per $j$ in $B$ and $5$ non-zero $l$’s per $k$ in $C$. In the latter, the first two loops iterate over the full matrix $C$ despite most of those iterations not leading to useful computation.
Formally, let $Q$ be the pointwise expression in our kernel, and let $Q_{(i_{1},\ldots,i_{k})}$ be the restriction of that expression to just the index variables $i_{1},\ldots,i_{k}$. Let $\mathbf{A}^{(i_{1},\ldots,i_{k})}$ be the input tensors that are not concordant with $i_{1},\ldots,i_{k}$. Then, we can define the cost of a loop order as follows,
@@
In practice, we further refine this model to take into account the number and kind of tensor accesses at each loop.
@@
To optimize the loop order, we combine this cost model with a branch-and-bound, dynamic programming algorithm. In the first pass, the optimization algorithm selects the cheapest loop index at each step until reaching a full loop order. This produces an upper bound on the optimal execution cost, which the algorithm uses to prune loop orders in the second step. This step applies a dynamic programming algorithm. Taking inspiration from Selinger’s algorithm for join ordering, each key in the DP table is a set of index variables and a set of inputs. The former represents the loops that have been iterated so far, and the latter represents a set of inputs that must be transposed.
@@ 5.2. Intermediate Formats
Once the loop order has been determined, the physical optimizer selects the optimal format for each query’s output. First, Galley sets the order of the indices to be concordant with either the loop order of the kernel where it will be consumed or the order requested by the user. Then, it selects a format for each index (e.g., dense vector, hash table, etc.). Two factors affect this decision: (1) the kind of writes being performed (sequential vs random) and (2) the sparsity of the tensor at this index. The former is important because many formats (e.g., sorted list formats) only allow sequential construction. These formats can only be used if the output indices form a prefix of the loop order.
When considering sparsity, Galley balances the fact that dense formats have better baseline efficiency, while sparse formats are asymptotically more efficient for highly sparse outputs. To describe this trade-off, we hand selected sparsity cutoffs between fully sparse, bytemap, and fully dense formats. To determine a particular output index’s format, the physical optimizer first determines the sparsity at this index level and uses our cutoffs to determine which category of formats to consider. Then, it checks whether sequential or random writes are being performed and selects the most efficient format that supports the write pattern.
@@ 6.1. Sparsity and Cardinality Estimation
Sparsity estimation is highly related to cardinality estimation in databases. However, translating methods for the latter to the former requires analyzing the algebraic properties of our tensor programs. For example, let $A_{ij}$ and $B_{jk}$ be sparse matrices with a fill value of 0, and let $R_{A}(I,J)$ and $R_{B}(J,K)$ be relations that store the indices of their non-zero entries. Assume we are performing the following,
@@
The correspondence results from the fact that 0 is the annihilator of multiplication (i.e., $x*0=0\forall x$), so any non-zero entry ijkijk in the output must correspond to a non-zero $ij$ in $A$ and a non-zero $jk$ in $B$. Consider the following instead:
@@
In this case, a nonzero ijkijk in the output can result from a non-zero $ij$ in $A$ or a non-zero $jk$ in $B$. In traditional relational algebra, where relations are over infinite domains, this kind of disjunction would result in an infinite relation. However, tensors have finite dimensions, so we can introduce relations that represent the finite domains of each index, e.g., $D_{i}=\{1,...,n_{i}\}$. This lets us represent the index relation of the output as
@@ 6.2. The Sparsity Statistics Interface
We use our statistics interface to annotate every node of the AST with statistics. Surprisingly, to support sparsity estimation over arbitrary tensor algebra expressions, we only need a few core functions: (1) a constructor, which produces statistics from a materialized tensor for Input and Alias nodes, (2-3) a function for (non) annihilating Map nodes (i.e., those whose children’s fill values are the annihilator of its pointwise function), which merges the children’s statistics, (4) a function for Agg, which adjusts the input’s statistics to reflect an aggregation over some set of indices, and (5) an estimation procedure, which estimates the sparsity of a tensor based on its statistics.
@@ 6.3.1. Uniform Estimator
The simplest statistic that can be kept about a tensor is the number of non-fill (e.g., non-zero) entries. The uniform estimator uses only this statistic and assumes these entries are uniformly distributed across the dimension space. This corresponds to System-R’s cardinality estimator with the added assumption that the active domain is the whole dimension for each index (Selinger et al., 1979).
Constructor. This function simply counts the non-fill values in the tensor, $nnz(A)$, and notes the dimension sizes $n_{i_{1}},\ldots,n_{i_{k}}$.
Map (Annihilating).
@@
Map (Non-Annihilating). To handle an non-annihilating pointwise operation, this function calculates the probability that an entry in the output was fill in all inputs. Then, it takes the compliment to get the probability that it was non-fill in all inputs and multiplies this by the output dimension space. Using the preceding notation:
@@
Aggregate. Given an input tensor $A_{I}$ to aggregate over the indices $I^{\prime}$, this function computes the probability that an output entry is non-fill by calculating the probability that at least one entry in the subspace of the input tensor was not fill:
@@
Estimate. The estimation function simply returns the current tensor’s stored cardinality statistic.
Example 0.
Suppose $A_{ij}$ and $B_{jk}$ are $100x100$ sparse matrices with $nnz(A)=1000$, $nnz(B)=200$, and we want to estimate $nnz(\sum_{j}A_{ij}B_{jk})$. We first compute $nnz(A_{ij}B_{jk})$ as $100^{3}*\frac{1000}{100^{2}}*\frac{200}{100^{2}}=2000$. The fractions are the probability that $i,j$ or $j,k$ was zero in $A$ or $B$, respectively. Next, we factor in the aggregation over $j$ to get $nnz(\sum_{j}A_{ij}B_{jk})=100^{2}*(1-(1-\frac{nnz(A_{ij}B_{jk})}{100^{3}})^{100})\approx 1800$. Here, the expression $(1-\frac{nnz(A_{ij}B_{jk})}{100^{3}})^{100}$ is the probability that all entries were zero for a particular $i,k$ pair.